CREAM Flight 2008
CREAM-IV at McMurdo, Nov. 25th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 25th of November at McMurdo.
- The CPT with the CDM completed today.
- The remaining 3 power phases (maximum power, battery mode, battery charging mode) are done.
- During the CPT, the instrument was in a flight mode. Everything went smoothly.
- All solar panels are assembled into the payload. A photo of the instrument with solar panels is uploaded.
CREAM-IV at McMurdo, Nov. 24th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 24th of November at McMurdo.
- The top SCD sparsification board Actel chips pins are cleaned again. Not much difference. But with the power is on, it seems improving in pedestal noise wise. Still some zero gain channels in the charge calibration run. Will send the muon run data files for checking the aliveness of sensors.
CREAM-IV at McMurdo, Nov. 23rd 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 23rd of November at McMurdo.
- CAL MB #7 temperature HSK display problem resolved by tightening the flat cable connector to the mother board side. After that, a small LPT for the calorimeter was done and OK.
- SFC cables, thermal fan-out box cables are cleaned-up for the CPT tomorrow.
CREAM-IV at McMurdo, Nov. 22nd 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 22nd of November at McMurdo.
- Weather briefing shows no significant changes in the stratosphere. Not so much winds for next few days.
- Parameter uploading in the Iridium mode test postponed to next time from WFF.
- Playback of remaining data set has completed.
CREAM-IV at McMurdo, Nov. 21st 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 21st of November at McMurdo.
- Weather briefing shows no significant changes in the stratosphere. Still the weather pattern analysis shows summer weather pattern to be 2 days later than the last year.
- All instruments were turned on except CAL HV from 8:30 AM for the End-to-End test in the Iridium connection mode.
- Data were taken in a flight mode successfully during the test.
- The reduced HSK data were received at SOC and displayed.



